a sample of 150 students were chosen at random from grade 11 students who have laptops.Pampanga. the sample indicated that 75% of them were in favor of having an educational trip. find a) 95% and b) 99% confidence intervals for the proportion if all students who are in favor having an educational trip


1
Expert's answer
2022-02-20T07:59:56-0500

a) "95\%CI=(0.75-1.96\sqrt{\frac{0.75(1-0.75)}{150}},0.75+1.96\sqrt{\frac{0.75(1-0.75)}{150}})=(0.68,0.82)."


b) "99\%CI=(0.75-2.58\sqrt{\frac{0.75(1-0.75)}{150}},0.75+2.58\sqrt{\frac{0.75(1-0.75)}{150}})=(0.66,0.84)."